OpenStudy (anonymous):

simplify by rationalizing the denominator \[\frac a{\sqrt{10}+\sqrt2}\]

OpenStudy (turingtest):

you should try to rewrite your question using the equation editor

OpenStudy (anonymous):

were is that?

OpenStudy (turingtest):

the \(\Sigma\) symbol button that says equation on the lower-left of the reply box

OpenStudy (anonymous):

oh ok thank u

OpenStudy (anonymous):

a\[\frac{ a }{ \sqrt{10+\sqrt{2}} }\]

OpenStudy (turingtest):

I'm guessing you meant\[\frac a{\sqrt{10}+\sqrt2}\]right?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

rationalizing means multiplying the numerator and denominator by the conjugate. the conjugate of (a+b) = (a-b)

OpenStudy (turingtest):

so multiply by\[\frac{\sqrt{10}-\sqrt2}{\sqrt{10}-\sqrt2}\]

OpenStudy (anonymous):

So that would cancel it out?

OpenStudy (turingtest):

well it is true that you can cancel the above to 1, but that's not going to do anything, is it? we multiply the numerator and denominator by something divided by itself precisely because it *is* equal to one, and therefor does not change the value of the expression.

OpenStudy (turingtest):

so just carry out the multiplication normally... remember what (a+b)(a-b) is....

OpenStudy (anonymous):

no

OpenStudy (turingtest):

well if you don't remember what (a+b)(a-b) equals just FOIL it out

OpenStudy (anonymous):

ok that is right

OpenStudy (anonymous):

\[a/\sqrt{20}\]

OpenStudy (turingtest):

no, we multiply top and bottom by \(\sqrt{10}-\sqrt2\) what is that times \(\sqrt{10}+\sqrt2\)

OpenStudy (turingtest):

?

OpenStudy (anonymous):

for some reason I am getting 10

OpenStudy (turingtest):

not quite... in general, what do you get by FOILing out (a+b)(a-b) ?

OpenStudy (anonymous):

you would times both a then b right?

OpenStudy (turingtest):

The FOIL method:if you don't remember how to FOIL you should go back a few steps and see what you missed in earlier algebra.